# TempoFormer: A Transformer for Temporally-aware Representations in Change Detection

### v1

@@ Abstract
Dynamic representation learning plays a pivotal role in understanding the evolution of linguistic content over time. On this front both context and time dynamics as well as their interplay are of prime importance. Current approaches model context via pre-trained representations, which are typically temporally agnostic. Previous work on modeling context and temporal dynamics has used recurrent methods, which are slow and prone to overfitting. Here we introduce TempoFormer, the fist task-agnostic transformer-based and temporally-aware model for dynamic representation learning. Our approach is jointly trained on inter and intra context dynamics and introduces a novel temporal variation of rotary positional embeddings. The architecture is flexible and can be used as the temporal representation foundation of other models or applied to different transformer-based architectures. We show new SOTA performance on three different real-time change detection tasks.
@@
Talia Tseriotou1, Adam Tsakalidis1,2,3 Maria Liakata1,2 1Queen Mary University of London, 2The Alan Turing Institute, 3European Centre for the Development of Vocational Training t.tseriotou@qmul.ac.uk
@@ 1 Introduction
Figure 1: The historical content provides important context towards identifying the Switch in the user’s mood from positive to negative in the last post.
Dynamic representation learning approaches aim to tackle this challenge. Dynamic word embedding methods have been studied in the context of semantic change detection (Bamler and Mandt, 2017; Rosenfeld and Erk, 2018). While changes in this context occur over long time periods, dynamic representation learning has been explored in other more temporally fine-grained tasks such as event detection (Yan et al., 2019; Yang et al., 2019; Lai et al., 2020), fake news detection (Vaibhav et al., 2019; Raza and Ding, 2022; Kaliyar et al., 2021) and mental health condition detection (Sawhney et al., 2021b; Tsakalidis et al., 2022a; Tseriotou et al., 2023). Nevertheless, dynamic representation learning research remains largely task or even dataset specific.
Transformer-based injection. The above mentioned approaches have relied on either pre-trained contextualised representations or transformer-based model layers (Devlin et al., 2018; Liu et al., 2019) to fine-tune representations before feeding them into RNN and CNN-like architectures as they had been shown to outperform transformer-based models (Ji et al., 2021; Gao et al., 2021; Tsakalidis et al., 2022b). However, since LSTM-based systems tend to overfit small datasets, transformer-based methods that overcome this issue would be a preferable method (Yu et al., 2020). Yet so far adapting layers on top of a transformer fails to strike the right balance between representation learning and task dynamics (Li et al., 2022; Ng et al., 2023).
@@
    Contrary to prior work reliant on pre-trained contextual embeddings, we allow for adaptation of transformers towards the domain and the temporal aspects of a dataset. We show that TempoFormer can be used as the foundation in more complex architectures (e.g. involving recurrence), striking the right balance between modeling a post/utterance (context-aware) and the timeline-level dynamics. Moreover the TempoFormer upper layers are flexible and can be applied in different Transformer-based architectures.
@@ 2 Related Work
Longitudinal Modeling and Change Detection: In addition to the importance of the linguistic stream, longitudinal tasks rely on temporal dynamics to asses the progression and identify changes over time. In the case of (a) identifying changes in user mood (Tsakalidis et al., 2022b; Hills et al., 2023; Tseriotou et al., 2023) and suicidal ideation through social media (Sawhney et al., 2021a) change is relative to the temporal evolution of users’ mood over time and approaches have relied mostly on recurrence and on utterance-level pretrained language model (PLM) representations. Tseriotou et al. (2024) introduced a longitudinal variation of (b) stance detection Yang et al. (2022); Kumar and Carley (2019) for detecting shifts (changes) in the public’s opinion towards an online rumour. They used Sentence-BERT (Reimers and Gurevych, 2019) representations with integration of path signatures (Lyons, 1998) in recurrence. For (c) conversation topic derailment, previous work has relied on fine-tuning transformer-based models (Konigari et al., 2021), providing extended context in their input (Kementchedjhieva and Søgaard, 2021) or applying recurrence over the utterance (Zhang et al., 2019a) and context stream (Chang and Danescu-Niculescu-Mizil, 2019). In this work we integrate stream dynamics directly into the transformer and show the flexibility of our approach as the foundation of different longitudinal tasks.
@@
Temporal Language Modeling: Many of the above tasks involve timestamps, which can enhance change detection through temporal dynamics. However, little research in NLP leverages time intervals and those who do assume equidistant time intervals between events (Ma and Gao, 2020; Tsakalidis and Liakata, 2020). Other work on temporal modelling has relied on hand crafted periodic task-specific time features (Kwon et al., 2013), concatenation of timestamp with linguistic representations (Tseriotou et al., 2023, 2024) or Hawkes temporal point process applied on top of recurrence (Guo et al., 2019; Hills et al., 2023). These approaches applied on top of LM representations miss the opportunity of training representations informed by temporal dynamics. Additionally, transformer-based models lack temporal sensitivity (Lazaridou et al., 2021; Loureiro et al., 2022). Rosin and Radinsky (2022) has conditioned attention weights on time, while Rosin et al. (2022); Wang et al. (2023) concatenated time tokens to text sequences. Although these methods create time-specific contextualised embeddings, they utilise absolute points in time rather than leveraging the temporal distance between units of textual information, important for context-aware and longitudinal tasks. Here we adapt the transformer attention mechanism to cater for the relative temporal aspect (§3.5).
@@
Hierarchical Models: Long content modeling approaches have leveraged transformer or attention-based blocks hierarchically on long documents, on input chunks/sentences and then on the sequence of such chunks (Zhang et al., 2019c; Pappagari et al., 2019; Wu et al., 2021; Li et al., 2023). This produces chunk-level summary embeddings, which preserve both the local and global aspects of contextualised representations. Here we leverage such local and global context dynamics to more efficiently model linguistic streams.
@@ 3.2 TempoFormer Overview
Fig. 2 provides an overview of TempoFormer. Its hierarchical architecture consists of three main modules, temporally-aware enhancement in multi-head attention to model the temporal distance between posts and a classification head. The modules are: post-level (local) encoding3.3) – obtaining word-level representation of each post using BERT’s first 10 layers; stream (global) encoding3.4) – modeling the sequential and temporal interactions between posts; and context-enhanced encoding3.6) – fusing stream-awareness in post-level representations to make them context-aware.
@@ 3.3 Post-level Encoding (Local)
Each training instance is a stream consisting of the current post and its recent history, alongside corresponding timestamps: [$[\{\mathbf{u}_{i-w+1},t_{i-w+1}\},...\{\mathbf{u}_{i-1},t_{i-1}\},\{\mathbf{u}_{i},t_{i}\}]$, with a total of $w$ posts in a stream. Timestamps are ignored at this stage. This stream of posts is converted into a stream, $e$, of word-level embeddings of word sequence length $K$ via the word and position embedding layer of BERT: [$[\{\mathbf{e}_{1,i-w+1},\mathbf{e}_{2,i-w+1}...,\mathbf{e}_{K,i-w+1}\},...$$\{\mathbf{e}_{1,i},\mathbf{e}_{2,i}...,\mathbf{e}_{K,i}\}]$. Specifically, in this module, the posts of each stream pass without post-post interactions via the first 10 BERT layers, resulting in hidden word-level representations for each post. Note that since a post is part of multiple streams through their window, it will pass through the BERT layers as part of each corresponding stream. For each post $j$ (belonging to a stream $q$), the word-level representations from the $z$-th Transformer layer are denoted as: $\mathbf{H}^{z}_{j_{q}}$ Therefore at the 10-th layer we reconstruct the stream and form local stream representation: $[\mathbf{H}^{10}_{i-w+1},...,\mathbf{H}^{10}_{i}]$.
@@ 3.5 Temporal Rotary Multi-Head Attention
Here, in order to model the temporal dynamics, we propose a novel variation of Eq. 1, named Temporal Rotary Multi-head Attention, making use of the relative position property. Instead of $R^{d}_{\theta,n-m}$, we reformulate the rotary matrix to model the temporal, rather than the positional differences, $R^{d}_{\theta,\mathbf{t_{n}-t_{m}}}$. We employ it on the stream-level using the [CLS] tokens to capture the stream global context through both the temporal and linguistic dynamics. The developed layer includes solely self-attention without the need for feed-forward and normalisation layers. In practice, since we measure time in seconds, we log-transform time in order to remove task dependancies on the scale of temporal propagation, to account for stream non-linearities and to alleviate exclusion of temporal outliers.
@@ 3.6 Context-enhanced Encoding
Literature has shown the effectiveness of enhancing word-level representations hierarchically through context-level learnt dynamics (Zheng et al., 2020; Wu et al., 2021; Ng et al., 2023). To this effect we introduce a second-layer of stream-level position embeddings, $\mathbf{s}^{11}$, to re-instate the absolute sequence position of each post for context-enhanced modeling. These are fed into a global context-aware layer, essentially a word-level transformer layer. Since the [CLS] tokens of each post are stream-aware, they contextualise the token-level representations based on the temporal and global learnt dynamics, obtaining: $[\mathbf{H}^{12}_{1,i-w+1},...,\mathbf{H}^{12}_{1,i}]$. To fully model the stream dynamics given the now context-enhanced [CLS] tokens, we employ a last layer of Temporal Rotary MHA resulting in $[\mathbf{H}^{{}^{\prime}12}_{1,i-w+1},...,\mathbf{H}^{{}^{\prime}12}_{1,i}]$. Lastly, we adapt the Gated Context Fusion (Gate&Norm) mechanism by Zheng et al. (2020) to fuse both the utterance word-level informed ($\mathbf{H}^{12}_{CLS}$) and the stream utterance-level informed ($\mathbf{H}^{{}^{\prime}12}_{CLS}$) [CLS] tokens through element-wise multiplication $\odot$:
@@ 3.7 Network Fine-Tuning
Although the proposed architecture can in principle be applied to any Transformer-based model, we select BERT (Devlin et al., 2018) as the foundation model and initialise all word-level weights. Literature on longitudinal context-aware classification has shown the importance of efficiently combining the current utterance with historical information (Sawhney et al., 2020, 2021a; Tseriotou et al., 2023). We thus concatenate the local stream-agnostic [CLS] token of the current utterance from the 10-th layer, $\mathbf{C}^{L}_{CLS}$, (obtained through typical BERT Pooling) with the obtained global stream-enhanced [CLS], $\mathbf{C}^{G}_{CLS}$ (Fig. 2). This final representation is fed through two fully connected layers with ReLU activation and dropout (Srivastava et al., 2014). The architecture is fine-tuned for each classification task (§4) using alpha-weighted focal loss (Lin et al., 2017), to assign more importance to minority classes and alleviate class imbalance.
@@ 4.1 Tasks and Datasets
We test our model on three different longitudinal tasks of different temporal granularity: 1) Stance Switch Detection – identification of switches in overall user stance around a social media claim, 2) Moments of Change (MoC) – identification of mood changes through users’ online posts and 3) Conversation Topic Shift – conversation diversion identification. We adopt a real-time prediction formulation (see §3.1) to assess system ability to perform early change detection in real-world scenarios. Table 1 provides the detailed statistics of each dataset showing the different degree of temporal granularity and dataset specifics. More details are provided in Appendix A.
Stance Switch Detection: Introduced by Tseriotou et al. (2024) this binary classification task tracks the support of posts towards a rumourous claim to identify switches in the overall stance. We use the LRS dataset based on RumourEval-2017 (Gorrell et al., 2019) and use N-Sw: to denote absence of switch and Sw: (switch) to denote a shift in the number of opposing vs supporting posts.
Moments of Change (MoC): Introduced by Tsakalidis et al. (2022b) this 3-class classification task assesses the mood changes of users given their online posting timelines. We use the TalkLife dataset with: IS: (switches) sudden mood shift from positive to negative (or v.v.); IE: (escalations) gradual mood progression from neutral or positive/negative to more positive/negative; and O: no mood change. In both datasets each post is accompanied by its timestamp.
Conversation Topic Shift: Given a corpus of conversations between humans, this binary classification task identifies utterance relevance to the main conversation topic versus another (derailment). We use the Topic Shift-MI (Mixed-Initiative) dataset (Konigari et al., 2021) annotated on a subset of Switch-board (Godfrey et al., 1992; Calhoun et al., 2010). This has a single but different major topic for each conversation, with M: (major) utterances belonging to the main topic and R: (rest) utterances that pertain to a different topic. Here conversations are not timestamped.

### v2

@@ Abstract
Dynamic representation learning plays a pivotal role in understanding the evolution of linguistic content over time. On this front both context and time dynamics as well as their interplay are of prime importance. Current approaches model context via pre-trained representations, which are typically temporally agnostic. Previous work on modelling context and temporal dynamics has used recurrent methods, which are slow and prone to overfitting. Here we introduce TempoFormer, the first task-agnostic transformer-based and temporally-aware model for dynamic representation learning. Our approach is jointly trained on inter and intra context dynamics and introduces a novel temporal variation of rotary positional embeddings. The architecture is flexible and can be used as the temporal representation foundation of other models or applied to different transformer-based architectures. We show new SOTA performance on three different real-time change detection tasks.
@@
Talia Tseriotou1, Adam Tsakalidis1,2 Maria Liakata1,2 1Queen Mary University of London, 2The Alan Turing Institute {t.tseriotou,m.liakata}@qmul.ac.uk
@@ 1 Introduction
Figure 1: Paraphrased example from the task of identifying moments of change in individuals’ mood (Tsakalidis et al., 2022b). Here, the historical content (light grey) provides important linguistic context towards identifying a Switch, a sudden mood shift from positive to negative, in the user’s behaviour at the last post (black).
Dynamic representation learning approaches aim to tackle this challenge. Dynamic word embedding methods have been studied in the context of semantic change detection (Bamler and Mandt, 2017; Rosenfeld and Erk, 2018). While changes in this context occur over long time periods, dynamic representation learning has been explored in other more temporally fine-grained tasks such as event detection (Yan et al., 2019; Yang et al., 2019; Lai et al., 2020), fake news detection (Vaibhav et al., 2019; Raza and Ding, 2022; Kaliyar et al., 2021) and mental health condition detection (Sawhney et al., 2021b; Tsakalidis et al., 2022a; Tseriotou et al., 2023). Such temporally fine-grained tasks significantly differ from semantic change detection approaches: not only on the temporal granularity aspect, but crucially with respect to event timeline length, irregularities in change frequency, annotation requirements and problem formulation. Therefore the adaptation of methodologies between the various sets of temporal change detection categories is at best challenging. Correspondingly, fine-grained dynamic representation learning research remains also largely task or even dataset specific.
Transformer-based injection. The above mentioned approaches have relied on either pre-trained contextualised representations or transformer-based model layers (Devlin et al., 2019; Liu et al., 2019) to fine-tune representations before feeding them into RNN and CNN-like architectures as so far they had been shown to outperform transformer-based models (Ji et al., 2021; Gao et al., 2021; Tsakalidis et al., 2022b). However, since LSTM-based systems tend to overfit small datasets, transformer-based methods that overcome this issue would be a preferable choice (Yu et al., 2020). Yet so far adapting layers on top of a transformer fails to strike the right balance between representation learning and task dynamics (Li et al., 2022; Ng et al., 2023).
@@
    Contrary to prior work reliant on pre-trained contextual embeddings, we allow for adaptation of transformers towards the domain and the temporal aspects of a dataset. We show that TempoFormer can be used as the foundation in more complex architectures (e.g. involving recurrence), striking the right balance between modelling a post/utterance (context-aware) and the timeline-level dynamics. Moreover the TempoFormer upper layers are flexible and can be applied in different Transformer-based architectures.
@@ 2 Related Work
Longitudinal Modelling and Change Detection: In addition to the importance of the linguistic stream, longitudinal tasks rely on temporal dynamics to asses progression and identify changes over time. In the case of (a) identifying changes in user mood (Tsakalidis et al., 2022b, a; Tseriotou et al., 2023; Hills et al., 2024) and suicidal ideation through social media (Sawhney et al., 2021a) change is relative to the temporal evolution of users’ mood over time and approaches have relied mostly on recurrence and on utterance-level pretrained language model (PLM) representations. Tseriotou et al. (2024) introduced a longitudinal variation of (b) stance detection Yang et al. (2022); Kumar and Carley (2019) for detecting shifts (changes) in the public opinion towards an online rumour. They used Sentence-BERT (Reimers and Gurevych, 2019) representations with integration of path signatures (Lyons, 1998) in recurrence. For (c) conversation topic derailment, previous work has relied on fine-tuning transformer-based models (Konigari et al., 2021), providing extended context in their input (Kementchedjhieva and Søgaard, 2021) or applying recurrence over the utterance (Zhang et al., 2019a) and context stream (Chang and Danescu-Niculescu-Mizil, 2019). In this work we integrate stream dynamics directly into the transformer and show the flexibility of our approach as the foundation of different longitudinal models.
@@
Temporal Language Modelling: Many of the above tasks involve timestamps, which can enhance change detection through temporal dynamics. However, little research in NLP leverages time intervals and those who do assume equidistant time intervals between events (Ma and Gao, 2020; Tsakalidis and Liakata, 2020). Other work on temporal modelling has relied on hand crafted periodic task-specific time features (Kwon et al., 2013), concatenation of timestamp with linguistic representations (Tseriotou et al., 2023, 2024) or Hawkes temporal point process applied on top of recurrence (Guo et al., 2019; Hills et al., 2024). These approaches applied on top of LM representations miss the opportunity of training representations informed by temporal dynamics. Additionally, transformer-based models lack temporal sensitivity (Lazaridou et al., 2021; Loureiro et al., 2022). Rosin and Radinsky (2022) has conditioned attention weights on time, while Rosin et al. (2022); Wang et al. (2023) concatenated time tokens to text sequences. Although these methods create time-specific contextualised embeddings, they utilise absolute points in time rather than leveraging the temporal distance between units of textual information, important for context-aware and longitudinal tasks. Here we adapt the transformer attention mechanism to cater for the relative temporal aspect (§3.5).
@@
Hierarchical Models: Long content modelling approaches have leveraged transformer or attention-based blocks hierarchically on long documents, on input chunks/sentences and then on the sequence of such chunks (Zhang et al., 2019c; Pappagari et al., 2019; Wu et al., 2021; Li et al., 2023). This produces chunk-level summary embeddings, which preserve both the local and global aspects of contextualised representations. Here we leverage such local and global context dynamics to more efficiently model linguistic streams.
@@ 3.2 TempoFormer Overview
Fig. 2 provides an overview of TempoFormer. Its hierarchical architecture consists of three main modules, temporally-aware enhancement in multi-head attention to model the temporal distance between posts and a classification head. The modules are: post-level (local) encoding3.3) – obtaining word-level representation of each post using BERT’s first 10 layers; stream (global) encoding3.4) – modelling the sequential and temporal interactions between posts; and context-enhanced encoding3.6) – fusing stream-awareness in post-level representations to make them context-aware.
@@ 3.3 Post-level Encoding (Local)
Each training instance is a stream consisting of the current post and its recent history, alongside corresponding timestamps: [$[\{\mathbf{u}_{i-w+1},t_{i-w+1}\},...\{\mathbf{u}_{i-1},t_{i-1}\},\{\mathbf{u}_{i},t_{i}\}]$, with a total of $w$ posts in a stream. Timestamps are ignored at this stage. This stream of posts is converted into a stream, $e$, of word-level embeddings of word sequence length $K$ via the word and position embedding layer of BERT: [$[\{\mathbf{e}_{1,i-w+1},\mathbf{e}_{2,i-w+1}...,\mathbf{e}_{K,i-w+1}\},...$$\{\mathbf{e}_{1,i},\mathbf{e}_{2,i}...,\mathbf{e}_{K,i}\}]$. Specifically, in this module, the posts in each stream pass without post-post interactions via the first 10 BERT layers, resulting in hidden word-level representations for each post. Note that since a post is part of multiple streams through their window, it will pass through the BERT layers as part of each corresponding stream. For each post $j$ (belonging to a stream $q$), the word-level representations from the $z$-th Transformer layer are denoted as: $\mathbf{H}^{z}_{j_{q}}$ Therefore at the 10-th layer we reconstruct the stream and form local stream representation: $[\mathbf{H}^{10}_{i-w+1},...,\mathbf{H}^{10}_{i}]$.
@@ 3.5 Temporal Rotary Multi-Head Attention
Here, in order to model the temporal dynamics, we propose a novel variation of Eq. 1, named Temporal Rotary Multi-head Attention, making use of the relative position property. Instead of $R^{d}_{\theta,n-m}$, we reformulate the rotary matrix to model the temporal, rather than the positional differences, $R^{d}_{\theta,\mathbf{t_{n}-t_{m}}}$. We employ it on the stream-level using the [CLS] tokens to capture the stream global context through both the temporal and linguistic dynamics. The developed layer includes solely self-attention without the need for feed-forward and normalisation layers. In practice, since we measure time in seconds, we log-transform time in order to remove task dependencies on the scale of temporal propagation, to account for stream non-linearities and to alleviate exclusion of temporal outliers.
@@ 3.6 Context-enhanced Encoding
Literature has shown the effectiveness of enhancing word-level representations hierarchically through context-level learnt dynamics (Zheng et al., 2021; Wu et al., 2021; Ng et al., 2023). To this effect we introduce a second-layer of stream-level position embeddings, $\mathbf{s}^{11}$, to re-instate the absolute sequence position of each post for context-enhanced modelling. These are fed into a global context-aware layer, essentially a word-level transformer layer. Since the [CLS] tokens of each post are stream-aware, they contextualise the token-level representations based on the temporal and global learnt dynamics, obtaining: $[\mathbf{H}^{12}_{1,i-w+1},...,\mathbf{H}^{12}_{1,i}]$. To fully model the stream dynamics given the now context-enhanced [CLS] tokens, we employ a last layer of Temporal Rotary MHA resulting in $[\mathbf{H}^{{}^{\prime}12}_{1,i-w+1},...,\mathbf{H}^{{}^{\prime}12}_{1,i}]$. Lastly, we adapt the Gated Context Fusion (Gate&Norm) mechanism by Zheng et al. (2021) to fuse both the utterance word-level informed ($\mathbf{H}^{12}_{CLS}$) and the stream utterance-level informed ($\mathbf{H}^{{}^{\prime}12}_{CLS}$) [CLS] tokens through element-wise multiplication $\odot$:
@@ 3.7 Network Fine-Tuning
Although the proposed architecture can in principle be applied to any Transformer-based model, we select BERT (Devlin et al., 2019) as the foundation model and initialise all word-level weights. Literature on longitudinal context-aware classification has shown the importance of efficiently combining the current utterance with historical information (Sawhney et al., 2020, 2021a; Tseriotou et al., 2023). We thus concatenate the local stream-agnostic [CLS] token of the current utterance from the 10-th layer, $\mathbf{C}^{L}_{CLS}$, (obtained through typical BERT Pooling) with the obtained global stream-enhanced [CLS], $\mathbf{C}^{G}_{CLS}$ (Fig. 2). This final representation is fed through two fully connected layers with ReLU activation and dropout (Srivastava et al., 2014). The architecture is fine-tuned for each classification task (§4) using alpha-weighted focal loss (Lin et al., 2017), to assign more importance to minority classes and alleviate class imbalance.
@@ 4.1 Tasks and Datasets
We test our model on three different longitudinal change detection classification tasks of different temporal granularity: 1) Stance Switch Detection – identification of switches in overall user stance around a social media claim, 2) Moments of Change (MoC) – identification of mood changes through users’ online posts and 3) Conversation Topic Shift – conversation diversion identification. We adopt a real-time prediction formulation (see §3.1) to assess system ability to perform early change detection in real-world scenarios. Table 1 provides detailed statistics for each dataset, showing the different degrees of temporal granularity and dataset specifics. More details on data splits and stream examples are provided in Appendix A and F respectively.
Stance Switch Detection: Introduced by Tseriotou et al. (2024) takes a sequence of chronologically ordered Twitter conversations about a rumourous claim related to a newsworthy event, to detect switches in the overall user stance. Conversations are converted from a tree structure into a chronologically ordered linear list (timeline). We use the LRS dataset from Tseriotou et al. (2024) based on RumourEval-2017 (Gorrell et al., 2019), and convert the original stance labels (supporting/denying/questioning/commenting) with respect to the root claim into two categories of Sw: (switch) a shift in the number of opposing (denying/questioning) vs supporting posts and N-Sw: absence of switch or cases where the numbers of supporting and opposing posts are equal. Each post is accompanied by its timestamp.
Moments of Change (MoC): Introduced by Tsakalidis et al. (2022b) takes a sequence of chronologically ordered posts shared by an online social media user, and classifies each post according to the behavioural change of the user as one of: IS- (switch) sudden mood shift from positive to negative (or vice versa); IE- (escalation) gradual mood progression from neutral or positive/negative to more positive/negative; and O- no mood change. We use the TalkLife dataset from Tsakalidis et al. (2022b) containing such annotated timelines where each post is timestamped.
Conversation Topic Shift: Given a corpus of open-domain conversations between humans, this binary classification task identifies whether each utterance falls under the main conversation topic or if it has derailed from it. We use the Topic Shift-MI (Mixed-Initiative) dataset (Konigari et al., 2021) annotated on a subset of the Switch-board dataset (Godfrey et al., 1992; Calhoun et al., 2010). This dataset has a single but varying major topic for each conversation. The two classes are M: (major) utterance belongs to the main topic and R: (rest) utterance pertains to a minor topic or is off-topic. Here conversations are not timestamped.